find omegen if we have 1500 revolutions per minute


1
Expert's answer
2021-02-24T12:49:16-0500
"\omega=1500\ \dfrac{rev}{min}\cdot\dfrac{1\ min}{60\ sec}\cdot2\pi\ rad=157\ \dfrac{rad}{s}."